# Mining Limited Data Sufficiently: A BERT-inspired Approach for CSI Time Series Application in Wireless Communication and Sensing

### v1

@@ III-B Data Process
Due to package loss, the received CSI series are always incomplete. Before inputting it to our model, we need to first identify where the package loss happens and use a placeholder [PAD] to occupy the place of lost packages. We judge whether package loss happens according to the gap between two successive timestamps. Given a sampling rate $f$, the ideal time gap between two successive timestamps should be $\Delta t=\frac{1}{f}$ seconds. Assume we receive a CSI series $[c_{1},c_{2},\dotsc,c_{m}]$ with corresponding timestamps $[t_{1},t_{2},\dotsc,t_{m}]$ with the unit of seconds. Then we can identify whether there is package loss and how many packages were lost between $c_{i}$ and $c_{i+1}$. We calculate the lost package amount according to:
$k=\text{round}\left(\frac{t_{i+1}-t_{i}}{\Delta t}\right)\ ,$ (2)
@@
Then we will add $k$ [PAD] tokens between $c_{i}$ and $c_{i+1}$. Besides, we also need to add $k$ corresponding timestamps between $t_{i}$ and $t_{i+1}$. We define the $j^{\text{th}}$ added timestamp between $t_{i}$ and $t_{i+1}$ as:
@@
where $\epsilon_{j}$ is a small noise. Our method to identify package loss is reasonable because after adding the [PAD] tokens, we found the package amount is very close to the theoretical amount when there is no package loss in our dataset.
@@ III-C Model Structure
The model structure of two generations of CSI-BERT is shown in Fig. 1. We utilize BERT [20] (the encoder part of a Transformer model [38]) as the backbone, but replace the bottom embedding layer and the top heads to adapt our data format and tasks.
III-C1 Embedding Layer
Compared to CSI-BERT1 [24], the main change is in the embedding layer. Our model’s embedding layer consists of three parts: token embedding, time embedding, and position embedding.
In the token embedding, we first use a standardization layer along the time dimension, as shown in Eq. 4:
$\mu^{(j)}=\frac{\sum_{i=1}^{n}c_{i}^{(j)}}{n}\ ,$ (4)
$\sigma^{(j)}=\sqrt{\frac{\sum_{i=1}^{n}(c_{i}^{(j)}-\mu^{(j)})^{2}}{n}}\ ,$
$\text{Standard}(c_{i}^{(j)})=\frac{c_{i}^{(j)}-\mu^{(j)}}{\sigma^{(j)}}\ ,$
where $c_{i}^{(j)}$ represents the $j^{th}$ dimension of $c_{i}$, $\mu$ and $\sigma$ represent the mean and standard deviation, respectively. We noticed that the distribution of CSI in each subcarrier changes significantly over time, so the standardization operation can help the model learn features better and mitigate the covariate shift problem [26, 39]. We also find it more efficient in CSI than other normalization operation like Batch Norm or Layer Norm.
Then, different from CSI-BERT1, we add an Adaptive Re-Weighting Layer (ARL) [40] along the subcarrier dimension to better capture the inner relationships, as shown in Eq. 5:
@@ III-C1 Embedding Layer
where $c$ is input, $T$ is output token, and MLP represents a Multilayer Perceptron. ARL is a mechanism similar to attention but simpler and more efficient. It first uses an MLP to generate a weight vector and then multiplies it with the original input. This process ensures that the network assigns higher significance to subcarriers with more informative content and ignores those without much variation, such as guard carriers. Thereby, it can help enhance the model’s precision and effectiveness. After that, another MLP will be used in the dimension of subcarrier to extract feature further.
For the time embedding layer, which accounts for the varying time intervals between consecutive CSI samples, we adopt a similar strategy as CSI-BERT1 but with an extra MLP to further embed the timestamp, as shown in Eq. 6:
$\text{TE}(t_{i})^{(j)}=\text{MLP}(\begin{matrix}\begin{cases}\text{sin}(\frac{\text{Norm}(t_{i})}{10^{\frac{4j}{d}}}),\ j=2k\\ \text{cos}(\frac{\text{Norm}(t_{i})}{10^{\frac{4(j-1)}{d}}}),\ j=2k+1\end{cases}\end{matrix})\ ,$ (6)
@@
where $j$ represents the $j^{th}$ dimension of $TE(t_{i})$, $k$ is a positive integer, $L$ is the maximum length of CSI sequence, and $d$ is the dimension of each CSI vector $c_{i}$. We first normalize the timestamp and then use the positional embedding method to encode it. However, due to the permutation-invariant nature of the Transformer architecture, it can only capture distance information but not order information. Studies [25] have also shown that Transformer-based models do not perform very well on time series data. To solve this problem, we use an MLP to capture time information further. Then, the time embedding is added to the token embedding and fused together using Layer Normalization.
Finally, the position embedding layer is the same as in the traditional Transformer model, and the position embedding result is added to the fusion of token embedding and time embedding. It is used to represent the relative position, while the time embedding is used to represent the absolute position.
III-C2 Head Layer
Similar to most BERT-based models, our CSI-BERT also has two branches for different tasks. The token-level head is used for CSI recovery and prediction tasks, while the sequence-level head is used for CSI classification tasks.
For the token-level head, a de-standardization layer is used to ensure that the output has a similar distribution as the input, as shown in Eq. 7:
@@ III-C2 Head Layer
where $y$ represents the output of the last linear layer, $\mu$ and $\sigma$ are calculated using Eq. 4. We consider the de-standardization layer to be necessary and efficient. First, because the distribution information is discarded by the standardization layer, we can only provide this information by the de-standardization layer. Then, we notice that the mean and standard deviation of CSI are relatively stable within a small time window. For example, in Fig. 4, we use the WiGesture dataset [WiGesture] to illustrate the average and standard deviation of CSI amplitude for each subcarrier in one second. We randomly select some packets from a total of 100 packets and show their average and standard deviation, corresponding to each line. The result shows that even with a change in the number of packets, the two statistics do not change significantly. It also means that within different packet loss rates, the two statistics in one second would not have a large difference. As a result, we consider the de-standardization operation can help the model obtain a relatively realistic distribution information no matter in recovery task or prediction task.
@@
Figure 4: Average and Standard Deviation of CSI Amplitude (WiGesture Dataset) [24]: Each line represents a different number of packets sampled from one second. The x-axis represents the subcarriers.
III-D Pre-training Phase
In the pre-training phase of CSI-BERT, we employ a similar approach to the MLM task in BERT [20]. This involves filling the empty positions with [PAD] tokens, randomly masking tokens with [MASK] tokens, and training the model to recover the masked tokens. However, due to the differences between continuous CSI and discrete natural language, we have made some modifications to adapt to the CSI data.
@@ III-D Pre-training Phase
In BERT, the [MASK] token is a fixed token, but in CSI-BERT, we replace the word embedding layer with a linear layer, which means the [MASK] token can have different values affecting the output. To address this, we assign a random value to the [MASK] token, sampled from a Gaussian distribution as shown in Eq. 8 whose mean and standard deviation are calculated from the standardization layer. This random approach makes it more challenging for the model to identify the position of the [MASK] token, thereby improving its understanding of the CSI sequence structure. Unlike CSI-BERT1, for the [PAD] token, even though we use a fixed value as [PAD], we replace it with [MASK] before inputting the CSI sequence to the model. This is because there is no [PAD] in the CSI recovery phase. Even though the [PAD] can be ignored by the attention mechanism through the attention mask matrix, it still affects the output of the ARL, which works in the dimension of subcarrier. If the [PAD] token only appears in the pre-training phase, it would cause a gap between inference and training, thereby influencing the model’s performance. As a result, we also cancel the attention matrix mechanism in this phase.
${}_{i}^{(j)}\sim\text{N}(\mu^{(j)},\sigma^{(j)})\ ,$ (8)
@@
For the pre-training phase, we follow a similar setting as RoBERTa [41] where some non-[PAD] tokens are randomly replaced with the [MASK] token. In RoBERTa, the masking proportion is fixed at 15%. However, in practical scenarios, the loss rate of CSI can vary, so we introduce a random masking proportion ranging from 15% to 70%, which is varied in each epoch.
Next, we train the model to recover these [MASK] tokens, where we refer to the model as the recoverer. To make the recovered CSI more realistic, we use an extra discriminator model. In CSI-BERT2, we use a GAN-based adversarial training strategy instead of the Domain Adversarial Neural Network (DANN) [42] based method in CSI-BERT1. While the DANN-based method has lower computational cost and faster training speed, the hyperparameter like gradient reversal rate $\lambda$ in it is not easy to set, which is always dynamic during training. We found the same $\lambda$ may not work when the training data changes, which may also cause the model to diverge. As a result, we select the Generative Adversarial Network (GAN) [43] based method in CSI-BERT2, where the discriminator does not share any parameters with the recoverer but has the same model structure. The optimization target of the GAN process can be represented as Eq. 9:
@@
where $D$ and $R$ represent the discriminator and recoverer, respectively, and $c$ represents the original CSI input. The discriminator tries to identify the original CSI input and the recovered CSI sequence output by the recoverer. Conversely, the recoverer tries to make it as confusing as possible, to make the recovered CSI more likely to be a real CSI sequence.
Specifically, the loss function of the recoverer consists of four parts, as shown in Eq. LABEL:Loss:
@@
where $\hat{c}$ represents the output of the recoverer, and $\hat{\mu}$ and $\hat{\sigma}$ represent the mean and standard deviation of $\hat{c}$ in the time dimension, respectively. The $L_{1}$ loss is the traditional loss function used in BERT, as it measures the accuracy of the output. However, in CSI recovery, the overall shape of the CSI is also important. Therefore, we utilize the $L_{2}$ and $L_{3}$ losses to consider the overall shape of the CSI by assessing its mean and standard deviation. Furthermore, the $L_{4}$ is the GAN loss for the recoverer, which tries to make the discriminator mistake in classifying the recovered CSI. Additionally, we calculate all the loss functions focusing only on the [MASK] tokens again to ensure that the model prioritizes the recovery of the missing CSI. What’s more, the loss function of the discriminator can be represented as Eq. 11:
@@
Besides, unlike CSI-BERT1 that divides the CSI by a sliding window before training, we randomly select a CSI sequence from the whole dataset during training. This is an approach to mitigate the risk of over-fitting, as the fixed data division may make the model dependent on the absolute position. However, the absolute position actually only depends on the data division.
III-E Fine-tuning Phase
III-E1 CSI Prediction Task
Encoder-based methods like BERT [20] are often considered not suitable enough for generation or prediction tasks due to their bi-directional attention mechanisms. This is also influenced by their pre-training task, where MLM is used to train the model to recover the inner data, not the future data. However, decoder-based methods like GPT [8] and encoder-decoder based methods like BART [7] also suffer from the problem of error accumulation [25]. Since they generate data in an autoregressive manner, an error may also influence the further generation results. In contrast, encoder-based methods have the advantage that they can generate all data in one step.
To make CSI-BERT2 more suitable for the prediction task, we design a Mask Prediction Model (MPM). MPM has a similar format as MLM, but it only masks tokens at the end of the sequence instead of masking tokens in arbitrary random positions. Specifically, we randomly mask 15% to 40% of the tokens at the end of the sequence. Except for this, the other training strategy is exactly the same as the pre-training.
III-E2 CSI Classification Task
For the CSI classification task, we adopt the approach commonly used for pre-trained models, where we freeze the bottom layers of CSI-BERT2 and only train the top layers for specific classification tasks. Different from CSI-BERT1, in this phase, we use [MASK] tokens to replace not only all [PAD] tokens but also some other input tokens randomly, to imitate the situation of different packet loss rates. By this way, our model has better generalization capacity, not limited to the scenarios in the training set.
III-F Inferring Phase
III-F1 CSI Prediction Task
In the inferring phase of the prediction task, we first use [MASK] to replace all [PAD] tokens, and then we fix the last 20% of tokens in the input sequences as [MASK]. The model is then used to predict the last 20% of tokens.
III-F2 CSI Classification Task
In the inferring phase of the classification task, there is no need to add extra [MASK] tokens as in fine-tuning, but only to use [MASK] to replace [PAD]. Furthermore, with the help of our time embedding, we find that our CSI-BERT2 model is able to effectively process data with different sampling rates, whereas traditional neural networks cannot always do this. For example, our model can successfully work when the testing data has a different sampling rate from the training data. The detailed experimental results will be shown in Section IV-C3.
@@ IV-A Experiment Configuration
We illustrate our model configurations in TableI. We trained our model on an NVIDIA RTX 3090. During training, we observed that CSI-BERT occupied approximately 2500MB of GPU memory.
@@ IV-B Dataset Description
In our experiment, we utilize three datasets: the publicly available WiGesture dataset [24], the WiFall Dataset [26], and a novel dataset named WiCount, which is proposed along with this paper. All these datasets are collected using the ESP32-S3 device as RX and a home router as TX, and share the same data format, with a sample rate of 100Hz, 1 antenna, and 52 subcarriers. For our experiment, we divide the data into 1-second samples, resulting in each sample having a length of 100 data points. We take the first 90% of the samples as the training set and the last 10% as the testing set within each category.

### v2

@@ III-B Data Pre-processing
Due to packet loss, the received CSI series are always incomplete. Before inputting them to our model, we need to first identify where the packet loss happens and use a placeholder [PAD] to occupy the place of lost packets. We determine the occurrence of packet loss by analyzing the time gap between two successive timestamps. Given a sampling rate $f$, the ideal time gap between two successive timestamps should be $\Delta t=\frac{1}{f}$ seconds. Assume we receive a CSI series $[c_{1},c_{2},\dotsc,c_{n}]$ with corresponding timestamps $[t_{1},t_{2},\dotsc,t_{n}]$ with the unit of seconds. Then we can identify whether there is packet loss and how many packets were lost between $c_{i}$ and $c_{i+1}$. We calculate the lost packet amount according to:
$k=\max\{\text{round}\left(\frac{t_{i+1}-t_{i}}{\Delta t}-1\right),0\}\ .$ (2)
@@
Then we will add $k$ [PAD] tokens between $c_{i}$ and $c_{i+1}$. Besides, we also need to add $k$ corresponding timestamps between $t_{i}$ and $t_{i+1}$. We define the $j-\text{th}$ added timestamp between $t_{i}$ and $t_{i+1}$ as:
@@
where $\epsilon_{j}$ is a small random number used to simulate the real-time gap. Our method to identify packet loss is reasonable because after adding the [PAD] tokens, we found the packet amount is very close to the theoretical amount when there is no packet loss in our dataset.
@@ III-C Model Structure
As shown in Fig. 1, we utilize BERT [18] as the backbone of CSI-BERT2, while replacing the bottom embedding layer and the top heads to accommodate our data format and tasks. The embedding layer consists of three components, which can be divided into two parts: token embedding for spatial encoding, and time embedding along with position embedding for temporal encoding. For the top heads, our design is based on the commonly used dual-branch head in Transformers, which includes a sequence-level head (where the entire sequence produces one output) and a token-level head (where each token generates one output).
III-C1 Spatial and Temporal Embedding Layer
To feed CSI data to the BERT backbone, we use an MLP-based method to extract features within each packet. However, in CSI data, different subcarriers contain varying amounts of information and importance. For example, guard carriers remain relatively consistent across packets, providing limited guidance for downstream tasks. Consequently, we aim for the model to learn the importance and relationships among subcarriers autonomously. The ARL module [20] is an effective choice, as it can adaptively assign different weights to different channels (subcarriers). The principle of ARL can be represented as:
@@ III-C1 Spatial and Temporal Embedding Layer
where $c$ is the input, and $T$ is the output token. ARL is a mechanism similar to attention but simpler and more efficient. It first employs an MLP to generate a weight vector and then multiplies this vector with the original input. This mechanism ensures that the network assigns greater importance to subcarriers containing informative content while down-weighting those with limited variation, such as guard carriers. As a result, it enhances the model’s precision and overall effectiveness. Subsequently, an additional MLP is applied along the subcarrier dimension to further extract features.
Additionally, prior to inputting the data, we follow CSI-BERT by using a standardization layer over the time (packet) dimension. This design is motivated by the observation that the distribution of CSI in each subcarrier changes significantly over time. Therefore, the time-dimension standardization operation can help the model learn useful features more effectively and increase its generalization capacity by mitigating the covariate shift problem [37, 38]. The operation can be represented as:
$\mu^{(j)}_{n}=\frac{\sum_{n=1}^{N}c_{n}^{(j)}}{N}\ ,$ (5)
$\sigma^{(j)}_{n}=\sqrt{\frac{\sum_{n=1}^{N}(c_{n}^{(j)}-\mu_{n}^{(j)})^{2}}{N}}\ ,$
$\text{Standard}(c_{n}^{(j)})=\frac{c_{n}^{(j)}-\mu_{n}^{(j)}}{\sigma_{n}^{(j)}}\ ,$
where $c_{n}^{(j)}$ represents the $j$-th dimension of $c_{n}$, and $\mu$ and $\sigma$ denote the mean and standard deviation, respectively.
Compared to the spatial encoding, which primarily focuses on the inner information within each packet, the temporal encoding aims to provide more order and position information to the BERT backbone for better learning of the relationships among packets. However, we noted that the vanilla positional embedding in Transformers is insufficient. First, the arrival gap between successive packets is not uniform due to various factors such as environmental noise and movement. The positional embedding cannot capture these small differences. Most importantly, it has been shown to be weak in time series data [19]. Due to the permutation-invariant nature of the Transformer architecture, it can only capture distance information but not order information.
To address these issues, we propose using an MLP to capture time information more effectively:
$\text{TE}(t_{i})=\text{MLP}(\text{PE}(t_{i}))\ ,$ (6)
[…]
@@
where TE and PE represent the time embedding and positional embedding, respectively, $\text{PE}(t_{i})^{(j)}$ denotes the $j$-th dimension of $\text{PE}(t_{i})$, $k$ is a positive integer, $L$ is the maximum length of the CSI sequence, and $d$ is the dimension of each CSI vector $c_{i}$. We first normalize the timestamp following the positional embedding method to encode it, and then feed it into the MLP for more complex information encoding. In this way, the original positional embedding captures the relative position information, while the time embedding encodes absolute temporal positions.
III-C2 Dual-branch Head Layer
CSI-BERT2 adopts a dual-branch architecture to support different tasks. The token-level head is used for CSI recovery and prediction tasks, while the sequence-level head is used for CSI classification tasks.
For the token-level head, a de-standardization layer is used to ensure that the output has a similar distribution as the input:
@@ III-C2 Dual-branch Head Layer
where $y$ represents the output of the last linear layer, $\mu$ and $\sigma$ are calculated using Eq. (5). The considered de-standardization layer is both necessary and efficient. This is because, on the one hand the distribution information is discarded by the standardization layer, we can only re-introduce this information by the de-standardization layer. On the other hand the mean and standard deviation of CSI are relatively stable within a small time window. For example, in Fig. 3, we use the WiGesture dataset [7] to illustrate the average and standard deviation of CSI amplitude for each subcarrier in one second. We randomly select a subset of packets from a total of 100, and show their average and standard deviation. The result shows that even with a change in the number of packets, the two statistics do not change significantly. It also implies that regardless of packet loss rate, these two statistics remain nearly consistent within a one-second interval. As a result, the de-standardization operation can help the model obtain a relatively realistic distribution information in both recovery task and prediction task.
@@
Figure 3: Average and standard deviation of CSI amplitude (WiGesture dataset [7]): Each line represents a different number of packets sampled from one second. The x-axis represents the subcarriers.
III-D Unsupervised Pre-training Process
In the pre-training phase of CSI-BERT2, we employ MLM to enable the model to learn basic data features in an unsupervised manner. Additionally, an adversarial learning strategy is also implemented to assist the model in recovering more realistic CSI data. Compared to CSI-BERT, we replace the adversarial learning approach of domain adversarial neural network (DANN) [40] with generative adversarial networks (GAN) [41]. This change is motivated by the observation that the performance of DANN heavily relies on the careful tuning of the gradient reversal coefficient $\lambda$, which can vary dynamically across different training conditions. Additionally, we found that a fixed $\lambda$ may not perform consistently when the training data changes, potentially causing the model to diverge. In contrast, the GAN-based approach offers greater stability. In the following sections, we provide the detailed introduction to the pre-training process.
@@ III-D Unsupervised Pre-training Process
For CSI sequence, the MLM involves filling empty positions with [PAD] tokens, randomly masking tokens with [MASK] tokens, and training the model to recover the masked tokens. In BERT, the value of [MASK] token is fixed, but in CSI-BERT2, we replace the word embedding layer with a linear layer, which means the [MASK] token can take on different values affecting the output. To address this, we assign a random value to [MASK] tokens, sampled from a Gaussian distribution $[MASK]_{i}^{(j)}\sim\text{N}(\mu_{i}^{(j)},\sigma_{i}^{(j)})$, where the mean and standard deviation are calculated from the standardization layer. This random approach makes it more challenging for the model to identify the positions of the [MASK] tokens, thereby encouraging a deeper understanding of the CSI sequence structure. For [PAD] token, even though we use a fixed value for [PAD], we replace them with [MASK] tokens before feeding the CSI sequence to the model. This is because there is no [PAD] in the CSI recovery phase. Even though the [PAD] can be ignored by the attention mechanism through the attention mask matrix, it still affects the output of the ARL, which works in the dimension of subcarrier. If [PAD] tokens only appear in the unsupervised training phase, it would cause a gap between inference and training, thereby influencing the model’s performance. As a result, we also disable the attention matrix mechanism in this phase.
@@
For the unsupervised training phase, we follow a similar setting as RoBERTa [42] where some non-[PAD] tokens are randomly replaced with [MASK] tokens. In RoBERTa, the masking proportion is fixed at 15%. However, in practical scenarios, the loss rate of CSI can vary, so we introduce a random masking proportion ranging from 15% to 70%, which is varied in each epoch.
Next, we train the model to recover the [MASK] tokens, referring to the model as the recoverer. To enhance the realism of the recovered CSI, we employ an additional discriminator model and train them using a GAN framework:
@@
where $D$ and $R$ represent the discriminator and recoverer, respectively, and $c$ represents the original CSI input. The discriminator tries to distinguish the original CSI input and the recovered CSI sequence output by the recoverer. The recoverer, in turn, learns to generate outputs that are indistinguishable from real CSI.
Specifically, the loss function of the recoverer consists of four parts:
@@
where $\hat{c}$ represents the output of the recoverer, and $\hat{\mu}$ and $\hat{\sigma}$ represent the mean and standard deviation of $\hat{c}$ in the time dimension, respectively. The $L_{1}$ loss is the traditional loss function used in BERT, as it measures the accuracy of the output. However, in CSI recovery, preserving the overall shape of the CSI is also important. Therefore, we utilize the $L_{2}$ and $L_{3}$ losses to consider the overall shape of the CSI by assessing its mean and standard deviation. Furthermore, the $L_{4}$ is the loss for the GAN training framework, which tries to make the discriminator mistake in distinguishing the recovered CSI. Additionally, we calculate all the loss functions focusing only on the [MASK] tokens again to ensure that the model prioritizes the recovery of the missing CSI. Moreover, the loss function of the discriminator is:
@@
To reduce overfitting, we also modify the training data selection strategy. Unlike CSI-BERT, which uses a sliding window, CSI-BERT2 randomly samples CSI sequences from the entire dataset during training.
III-E Fine-tuning and Inference Process in Downstream Tasks
III-E1 CSI Prediction
Encoder-based methods like BERT [18] are often considered not suitable enough for generation or prediction tasks due to their bi-directional attention mechanisms. This is also influenced by their unsupervised training task, where MLM is used to train the model to recover the inner data, not the future data. However, decoder-based methods like GPT [43] and encoder-decoder based methods like BART [44] also suffer from the problem of error accumulation [19]. Since they generate data in an autoregressive manner, an error may also influence the further generation results. In contrast, encoder-based methods have the advantage that they can generate all data in one step.
To overcome this, CSI-BERT2 introduces a MPM training method. MPM has a similar format as MLM, but it only masks tokens at the end of the sequence instead of masking tokens in arbitrary random positions. Specifically, we randomly mask 15% to 40% of the tokens at the end of the sequence. Except for this, the other training strategy is exactly the same as the unsupervised training.
In the inference phase of the prediction task, we first use [MASK] to replace all [PAD] tokens, and then we fix the last 20% of tokens in the input sequences as [MASK]. The model is then used to predict the last 20% of tokens.
III-E2 CSI Classification
For CSI classification tasks, we adopt the approach commonly used for pre-trained models, where we freeze the bottom layers of CSI-BERT2 and only train the top layers for task-specific adaptation. Different from CSI-BERT, in this phase, we use [MASK] tokens to replace not only all [PAD] tokens but also some other input tokens randomly, to imitate the situation of different packet loss rates. This improves generalization to real-world scenarios where packet loss rates may vary during deployment.
In the inference phase of the classification task, there is no need to add extra [MASK] tokens as in fine-tuning, but only to use [MASK] to replace [PAD]. Furthermore, thanks to our time embedding, CSI-BERT2 generalizes effectively to sequences with different sampling rates. Unlike traditional neural networks, which often fail under such variability, CSI-BERT2 maintains high performance even when the sampling rate of the testing data differs from that used in training.
@@ IV-A Experiment Configuration
We illustrate our model configurations in Table I. We trained our model on an NVIDIA RTX 3090 GPU. During training, we observed that CSI-BERT2 occupies approximately 2500MB of GPU memory.
@@ IV-B Dataset Description
In our experiment, we utilize three real-collected Wi-Fi sensing datasets: the publicly available WiGesture dataset [7], the WiFall Dataset [37], and a novel dataset named WiCount, which is proposed along with this paper. All these datasets are collected using the ESP32-S3 device as RX and a home router as TX, and share the same data format, with a sample rate of 100Hz, 1 antenna, and 52 subcarriers. Additionally, considering the communication scenario, due to the limited availability of open-source public datasets, we generate a simulated dataset named CommPre using Sionna [45], which maintains the same format as other three sensing datasets. In our experiment, we divide the data into 1-second samples, resulting in each sample having a length of 100 data points. We take the first 90% of the samples as the training set and the remaining 10% as the testing set.